# Integrating large language models and active inference to understand eye movements in reading and dyslexia

### v2

@@ Acknowledgments
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreements No. 945539 (Human Brain Project SGA3) and No. 952215 (TAILOR) to GP; the European Research Council under the Grant Agreement No. 820213 (ThinkAhead) to GP; the PNRR MUR projects PE0000013-FAIR and IR0000011–EBRAINS-Italy to GP; the MUR PRIN2020 project “Free energy principle and the brain: neuronal and phylogenetic mechanisms of Bayesian inference” - Grant N. 2020529PCP to FD. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@
Francesco Donnarumma
Mirco Frosolone
Giovanni Pezzulo
@@
Via San Martino della Battaglia, 44
@@
4 - letter 8 - letter
@@
Control Model $100$ $1.00\pm 1.03\cdot 10^{-14}$ $100$ $1.00\pm 7.85\cdot 10^{-15}$
DM -Noise on Level 1 $100$ $1.00\pm 3.32\cdot 10^{-04}$ $100$ $0.99\pm 3.82\cdot 10^{-04}$
@@
4 - letter 8 - letter
@@
4 - letter 8 - letter
@@
4 - letter 8 - letter
@@
4 - letter 8 - letter
@@
4 - word 8 - word
@@
Control Model $100$ $1.00\pm 3.66\cdot 10^{-16}$ $100$ $1.00\pm 3.21\cdot 10^{-12}$
DM - Noise on Level 1 $100$ $1.00\pm 3.64\cdot 10^{-06}$ $100$ $1.00\pm 5.69\cdot 10^{-08}$
DM - Noise on Level 2 $100$ $1.00\pm 3.40\cdot 10^{-07}$ $100$ $1.00\pm 3.21\cdot 10^{-12}$
@@
4 - word 8 - word
DM - Noise Level 1 $F[1,198]=0.02,~{}p\approx 0.90$ $F[1,198]=0.01,~{}p\approx 0.92$
DM - Noise Level 2 $F[1,198]=0.97,~{}p\approx 0.453$ $F[1,198]=0.451,~{}~{}p\approx 0.58$
@@
4 - word 8 - word
DM - Noise Level 1 $F[1,198]=0.451,~{}p\approx 0.52$ $F[1,198]=0.82,~{}p\approx 0.457$
DM - Noise Level 2 $F[1,198]=0.01,~{}p\approx 0.90$ $F[1,198]=1.86,~{}p\approx 0.17$
@@
4 - word 8 - word
@@
4 - word 8 - word
@@
Flat Priors VS Informative Priors $F[1,198]=4.07,\mathbf{{\bf\it p}\approx 4.1\cdot 10^{-03}}$
@@
Flat Priors VS Informative Priors $F[1,198]=2.57,~{}~{}p\approx 0.11$
@@
Simulation 2 (4 - word) $25.96$ $1.96$ $21$ $31$
Simulation 2 (8 - word) $50.24$ $1.64$ $47$ $55$
@@ IX MPH LANG LOVER LILITH IRANIAN GROUPING
RR SHU STYX SQUAT SAXONY REISSUE SHOWERED
@@ XI YET WINS WRAPS WASHED VACANCY WARRIORS
FAITH CONQUERS ALL OBSTACLES
FAITH CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH CONQUERS EACH HURDLE
FAITH CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH CONQUERS EACH SETBACK
FAITH CONQUERS EACH HURDLE BUT TRASCENDS ALL MISFORTUNE
FAITH CONQUERS EVERY CHALLENGES
@@
FAITH CONQUERS EACH SETBACK BUT SURMOUNTS ALL OBSTACLES
FAITH CONQUERS EVERY OBSTACLES
FAITH CONQUERS EVERY OBSTACLES AND PREVAILS OVER ADVERSITY
FAITH DEFEATS ALL OBSTACLES
FAITH CONQUERS EVERY OBSTACLES AND SURMOUNTS ALL OBSTACLES
FAITH OVERCOMES ALL HINDRANCES
FAITH DEFEATS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH OVERCOMES ALL OBSTACLES
FAITH DEFEATS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH OVERPOWERS ALL LIMITATIONS
@@
FAITH OVERCOMES ALL HINDRANCES AND TRASCENDS OVER MISFORTUNE
FAITH PREVAILS AGAINST BARRIERS
@@
FAITH OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH PREVAILS AGAINST OBSTACLES
FAITH OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
FAITH PREVAILS OVER EVERY ADVERSITY
FAITH OVERPOWERS ALL LIMITATIONS BUT TRASCENDS ALL MISFORTUNE
FAITH PREVAILS OVER OBSTACLES
@@
FAITH PREVAILS AGAINST BARRIERS AND TRASCENDS OVER MISFORTUNE
FAITH SURMOUNTS ALL OBSTACLES
FAITH PREVAILS AGAINST OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH SURPASSES ALL OBSTACLES
FAITH PREVAILS EVERY ADVERSITY BUT SURMOUNTS ALL OBSTACLES
FAITH SURPASSES ALL OPPOSITION
@@
FAITH PREVAILS OVER ADVERSITY BUT SURMOUNTS ALL OBSTACLES
FAITH TRIUMPHS AGAINST OBSTACLES
@@
FAITH PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH TRIUMPHS OVER CHALLENGES
FAITH SURMOUNTS ALL OBSTACLES AND SURMOUNTS ALL MISFORTUNE
FAITH TRIUMPHS OVER OBSTACLES
FAITH SURPASSES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH VANQUISHES ALL OBSTACLES
FAITH SURPASSES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH VANQUISHES EVERY DIFFICULTY
@@
FAITH SURPASSES ALL OBSTACLES BUT SURMOUNTS ALL ADVERSITY
HOPE CONQUERS ALL LIMITATIONS
@@
FAITH SURPASSES ALL OPPOSITION BUT SURMOUNTS ALL MISFORTUNE
HOPE CONQUERS ALL OBSTACLES
@@
FAITH TRIUMPHS AGAINST OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH CHALLENGES
@@
FAITH TRIUMPHS OVER CHALLENGES AND TRASCENDS OVER MISFORTUNE
HOPE CONQUERS EACH HINDRANCES
@@
FAITH TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EACH HURDLE
@@
FAITH TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH INHIBITION
@@
FAITH VANQUISHES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH LIMITATIONS
@@
FAITH VANQUISHES EVERY DIFFICULTY BUT TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH SETBACK
HOPE CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY ADVERSITY
HOPE CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EVERY BARRIERS
@@
HOPE CONQUERS EACH SETBACK BUT SURMOUNTS ALL OBSTACLES
HOPE CONQUERS EVERY CHALLENGES
@@
HOPE DEFEATS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY DIFFICULTY
@@
HOPE DEFEATS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EVERY HURDLE
@@
HOPE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY OBSTACLES
HOPE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EVERY RESTRAINT
@@
HOPE PREVAILS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE DEFEATS ALL OPPOSITION
@@
HOPE PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE OVERCOMES ALL CHALLENGES
HOPE SURPASSES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE OVERCOMES EACH BOUNDARY
@@
HOPE SURPASSES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE OVERCOMES EACH HINDRANCES
@@
HOPE TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE PREVAILS AGAINST ADVERSITY
@@
HOPE TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE PREVAILS AGAINST BARRIERS
LOVE CONQUERS ALL ADVERSITY AND TRASCENDS ALL ADVERSITY
HOPE PREVAILS AGAINST IMPEDIMENT
@@
LOVE CONQUERS ALL ADVERSITY AND TRASCENDS ALL MISFORTUNE
HOPE PREVAILS OVER OBSTACLES
@@
LOVE CONQUERS ALL ADVERSITY BUT TRASCENDS ALL MISFORTUNE
HOPE SURMOUNTS ALL DIFFICULTIES
@@
LOVE CONQUERS ALL ADVERSITY BUT TRASCENDS EVERY MISFORTUNE
HOPE SURPASSES ALL OBSTACLES
@@
LOVE CONQUERS ALL BOUNDARIES AND TRASCENDS ALL ADVERSITY
HOPE SURPASSES EVERY CONFINEMENT
LOVE CONQUERS ALL BOUNDARIES AND TRASCENDS ALL MISFORTUNE
HOPE SURPASSES EVERY OBSTACLES
LOVE CONQUERS ALL BOUNDARIES BUT TRASCENDS ALL MISFORTUNE
HOPE TRANSCENDS EACH RESTRICTION
@@
LOVE CONQUERS ALL BOUNDARIES BUT TRASCENDS EVERY MISFORTUNE
HOPE TRIUMPHS AGAINST OBSTACLES
LOVE CONQUERS ALL DIFFERENCES AND TRASCENDS ALL ADVERSITY
HOPE TRIUMPHS OVER CHALLENGES
LOVE CONQUERS ALL DIFFERENCES AND TRASCENDS ALL MISFORTUNE
HOPE VANQUISHES ALL HINDRANCES
LOVE CONQUERS ALL DIFFERENCES BUT TRASCENDS ALL MISFORTUNE
LOVE OVERCOMES EACH HINDRANCES
LOVE CONQUERS ALL DIFFERENCES BUT TRASCENDS EVERY MISFORTUNE
LOVE CONQUERS ALL ADVERSITY
LOVE CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
LOVE CONQUERS ALL BOUNDARIES
LOVE CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
LOVE CONQUERS ALL DIFFERENCES
LOVE CONQUERS ALL OBSTACLES BUT TRASCENDS ALL ADVERSITY
LOVE CONQUERS ALL OBSTACLES
LOVE CONQUERS ALL OBSTACLES BUT TRASCENDS EVERY ADVERSITY
LOVE CONQUERS ANY OBSTACLES
LOVE CONQUERS EVERY BARRIERS AND PREVAILS OVER ADVERSITY
LOVE CONQUERS EACH OBSTACLES
@@
LOVE CONQUERS EVERY OBSTACLES BUT TRASCENDS ALL MISFORTUNE
LOVE CONQUERS EVERY BARRIERS
LOVE DEFEATS ALL OBSTACLES AND PREVAILS OVER ADVERSITY
LOVE CONQUERS EVERY CHALLENGES
LOVE OVERCOMES ALL OBSTACLES AND CONQUERS ALL ADVERSITY
LOVE CONQUERS EVERY CONFINEMENT
LOVE OVERCOMES ALL OBSTACLES AND CONQUERS ALL MISFORTUNE
LOVE CONQUERS EVERY DIFFICULTY
LOVE OVERCOMES ALL OBSTACLES AND CONQUERS EVERY OBSTACLES
LOVE CONQUERS EVERY HINDRANCES
LOVE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
LOVE CONQUERS EVERY HURDLE
LOVE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
LOVE CONQUERS EVERY LIMITATIONS
LOVE OVERCOMES ALL OBSTACLES BUT CONQUERS ALL ADVERSITY
LOVE CONQUERS EVERY OBSTACLES
LOVE OVERCOMES ALL OBSTACLES BUT CONQUERS EVERY ADVERSITY
LOVE CONQUERS EVERY SETBACK
LOVE OVERCOMES ALL OBSTACLES BUT TRASCENDS ALL ADVERSITY
LOVE DEFEATS ALL OBSTACLES
LOVE OVERCOMES ALL OBSTACLES BUT TRASCENDS EACH ADVERSITY
LOVE DEFEATS ALL OPPOSITION
LOVE OVERCOMES ALL OBSTACLES BUT TRASCENDS EVERY ADVERSITY
LOVE OVERCOMES ALL OBSTACLES
LOVE PREVAILS AGAINST CHALLENGES AND PREVAILS OVER ADVERSITY
LOVE PREVAILS AGAINST CHALLENGES
LOVE PREVAILS AGAINST OBSTACLES AND OVERCOMES OVER MISFORTUNE
LOVE PREVAILS AGAINST OBSTACLES
LOVE PREVAILS AGAINST OBSTACLES AND PREVAILS OVER ADVERSITY
LOVE PREVAILS OVER OBSTACLES
LOVE PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
LOVE SURMOUNTS ALL OBSTACLES
LOVE SURMOUNTS ALL OBSTACLES BUT TRASCENDS ALL ADVERSITY
LOVE SURPASSES ALL OBSTACLES
LOVE SURMOUNTS ALL OBSTACLES BUT TRASCENDS ALL MISFORTUNE
LOVE TRIUMPHS AGAINST OBSTACLES
LOVE SURPASSES ALL OBSTACLES AND PREVAILS OVER ADVERSITY
LOVE TRIUMPHS OVER OBSTACLES
@@
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS ALL ADVERSITY
LOVE VANQUISHES ALL OBSTACLES
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS ALL OBSTACLES
UNITY BREAKS THROUGH ALL BARRIERS
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS EACH ADVERSITY
UNITY CONQUERS AGAINST ODDS
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS OVER ADVERSITY
UNITY CONQUERS ALL OBSTACLES
LOVE TRIUMPHS AGAINST OBSTACLES AND DEFEATS OVER OBSTACLES
UNITY CONQUERS DESPITE ODDS
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS ALL ADVERSITY
UNITY DEFEATS AGAINST ODDS
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS ALL OBSTACLES
UNITY DEFEATS ALL OPPOSITION
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS EACH ADVERSITY
UNITY OVERCOMES AGAINST ODDS
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS OVER ADVERSITY
UNITY OVERCOMES ALL CHALLENGES
LOVE TRIUMPHS AGAINST OBSTACLES BUT DEFEATS OVER OBSTACLES
UNITY OVERPOWERS ALL RESISTANCE

### v3

@@ Acknowledgments
This research received funding from the European Research Council under the Grant Agreement No. 820213 (ThinkAhead), the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”; Project PE0000013, “FAIR”; Project PE0000006, “MNESYS”), and the Ministry of University and Research, PRIN PNRR P20224FESY, PRIN 20229Z7M8N and PRIN 2020529PCP. The GEFORCE Quadro RTX6000 and Titan GPU cards used for this research were donated by the NVIDIA Corporation. We used a Generative AI model to correct typographical errors and edit language for clarity.
@@
Francesco Donnarumma Mirco Frosolone Giovanni Pezzulo
@@
Via Gian Domenico Romagnosi, 18A
@@
4 - letter word 8 - letter word
@@
Control Model $100$ $\pm 1.03\cdot 10^{-14} 100100$ $1.00\pm 7.85\cdot 10^{-15}$
DM -Noise on Level 1 $100$ $\pm 3.32\cdot 10^{-04} 100100$ $0.99\pm 3.82\cdot 10^{-04}$
@@
4 - letter word 8 - letter word
@@
4 - letter word 8 - letter word
@@
4 - letter word 8 - letter word
@@
4 - letter word 8 - letter word
@@
4 - word sentence 8 - word sentence
@@
Control Model $100$ $\pm 3.66\cdot 10^{-16} 100100$ $1.00\pm 3.21\cdot 10^{-12}$
DM - Noise on Level 1 $100$ $\pm 3.64\cdot 10^{-06} 100100$ $1.00\pm 5.69\cdot 10^{-08}$
DM - Noise on Level 2 $100$ $\pm 3.40\cdot 10^{-07} 100100$ $1.00\pm 3.21\cdot 10^{-12}$
@@
4 - word sentence 8 - word sentence
DM - Noise Level 1 $F[1,198]=0.02,\leavevmode\nobreak\ p\approx 0.90$ $F[1,198]=0.01,\leavevmode\nobreak\ p\approx 0.92$
DM - Noise Level 2 $F[1,198]=0.97,\leavevmode\nobreak\ p\approx 0.45$ $F[1,198]=0.45,\leavevmode\nobreak\ \leavevmode\nobreak\ p\approx 0.58$
@@
4 - word sentence 8 - word sentence
DM - Noise Level 1 $F[1,198]=0.45,\leavevmode\nobreak\ p\approx 0.52$ $F[1,198]=0.82,\leavevmode\nobreak\ p\approx 0.46$
DM - Noise Level 2 $F[1,198]=0.01,\leavevmode\nobreak\ p\approx 0.90$ $F[1,198]=1.86,\leavevmode\nobreak\ p\approx 0.17$
@@
4 - word sentence 8 - word sentence
@@
4 - word sentence 8 - word sentence
@@
Flat Priors VS Informative Priors $F[1,198]=4.1,\mathbf{{\bf\it p}\approx 4.1\cdot 10^{-03}}$
@@
Flat Priors VS Informative Priors $F[1,198]=2.6,\leavevmode\nobreak\ \leavevmode\nobreak\ p\approx 0.11$
@@
Simulation 2 (4 words) $25.96$ $1.96$ 2121 $31$
Simulation 2 (8 words) $50.24$ $1.64$ $47$ $55$
Simulation 3 (4 words - 1 Unknown) $25.46$ $1.92$ $22$ $32$
@@ IX MPH LANG LOVER LILITH IRANIAN GROUPING
$R$ SHU STYX SQUAT SAXONY REISSUE SHOWERED
@@ XI YET WINS WRAPS WASHED VACANCY WARRIORS
FAITH CONQUERS ALL OBSTACLES FAITH CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH CONQUERS EACH HURDLE FAITH CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH CONQUERS EACH SETBACK FAITH CONQUERS EACH HURDLE BUT TRASCENDS ALL MISFORTUNE
@@
FAITH CONQUERS EVERY CHALLENGES FAITH CONQUERS EACH SETBACK BUT SURMOUNTS ALL OBSTACLES
FAITH CONQUERS EVERY OBSTACLES FAITH CONQUERS EVERY OBSTACLES AND PREVAILS OVER ADVERSITY
FAITH DEFEATS ALL OBSTACLES FAITH CONQUERS EVERY OBSTACLES AND SURMOUNTS ALL OBSTACLES
FAITH OVERCOMES ALL HINDRANCES FAITH DEFEATS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH OVERCOMES ALL OBSTACLES FAITH DEFEATS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
FAITH OVERPOWERS ALL LIMITATIONS FAITH OVERCOMES ALL HINDRANCES AND TRASCENDS OVER MISFORTUNE
@@
FAITH PREVAILS AGAINST BARRIERS FAITH OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH PREVAILS AGAINST OBSTACLES FAITH OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
FAITH PREVAILS OVER EVERY ADVERSITY FAITH OVERPOWERS ALL LIMITATIONS BUT TRASCENDS ALL MISFORTUNE
@@
FAITH PREVAILS OVER OBSTACLES FAITH PREVAILS AGAINST BARRIERS AND TRASCENDS OVER MISFORTUNE
FAITH SURMOUNTS ALL OBSTACLES FAITH PREVAILS AGAINST OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH SURPASSES ALL OBSTACLES FAITH PREVAILS EVERY ADVERSITY BUT SURMOUNTS ALL OBSTACLES
@@
FAITH SURPASSES ALL OPPOSITION FAITH PREVAILS OVER ADVERSITY BUT SURMOUNTS ALL OBSTACLES
@@
FAITH TRIUMPHS AGAINST OBSTACLES FAITH PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
FAITH TRIUMPHS OVER CHALLENGES FAITH SURMOUNTS ALL OBSTACLES AND SURMOUNTS ALL MISFORTUNE
FAITH TRIUMPHS OVER OBSTACLES FAITH SURPASSES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
FAITH VANQUISHES ALL OBSTACLES FAITH SURPASSES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
FAITH VANQUISHES EVERY DIFFICULTY FAITH SURPASSES ALL OBSTACLES BUT SURMOUNTS ALL ADVERSITY
@@
HOPE CONQUERS ALL LIMITATIONS FAITH SURPASSES ALL OPPOSITION BUT SURMOUNTS ALL MISFORTUNE
@@
HOPE CONQUERS ALL OBSTACLES FAITH TRIUMPHS AGAINST OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE CONQUERS EACH CHALLENGES FAITH TRIUMPHS OVER CHALLENGES AND TRASCENDS OVER MISFORTUNE
@@
HOPE CONQUERS EACH HINDRANCES FAITH TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
@@
HOPE CONQUERS EACH HURDLE FAITH TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE CONQUERS EACH INHIBITION FAITH VANQUISHES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE CONQUERS EACH LIMITATIONS FAITH VANQUISHES EVERY DIFFICULTY BUT TRASCENDS ALL MISFORTUNE
HOPE CONQUERS EACH SETBACK HOPE CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY ADVERSITY HOPE CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE CONQUERS EVERY BARRIERS HOPE CONQUERS EACH SETBACK BUT SURMOUNTS ALL OBSTACLES
@@
HOPE CONQUERS EVERY CHALLENGES HOPE DEFEATS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
@@
HOPE CONQUERS EVERY DIFFICULTY HOPE DEFEATS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE CONQUERS EVERY HURDLE HOPE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
HOPE CONQUERS EVERY OBSTACLES HOPE OVERCOMES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE CONQUERS EVERY RESTRAINT HOPE PREVAILS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
@@
HOPE DEFEATS ALL OPPOSITION HOPE PREVAILS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
HOPE OVERCOMES ALL CHALLENGES HOPE SURPASSES ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
@@
HOPE OVERCOMES EACH BOUNDARY HOPE SURPASSES ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE OVERCOMES EACH HINDRANCES HOPE TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL ADVERSITY
@@
HOPE PREVAILS AGAINST ADVERSITY HOPE TRIUMPHS OVER OBSTACLES AND TRASCENDS ALL MISFORTUNE
@@
HOPE PREVAILS AGAINST BARRIERS LOVE CONQUERS ALL ADVERSITY AND TRASCENDS ALL ADVERSITY
@@
HOPE PREVAILS AGAINST IMPEDIMENT LOVE CONQUERS ALL ADVERSITY AND TRASCENDS ALL MISFORTUNE
@@
HOPE PREVAILS OVER OBSTACLES LOVE CONQUERS ALL ADVERSITY BUT TRASCENDS ALL MISFORTUNE
@@
HOPE SURMOUNTS ALL DIFFICULTIES LOVE CONQUERS ALL ADVERSITY BUT TRASCENDS EVERY MISFORTUNE
@@
HOPE SURPASSES ALL OBSTACLES LOVE CONQUERS ALL BOUNDARIES AND TRASCENDS ALL ADVERSITY
@@
HOPE SURPASSES EVERY CONFINEMENT LOVE CONQUERS ALL BOUNDARIES AND TRASCENDS ALL MISFORTUNE
@@
HOPE SURPASSES EVERY OBSTACLES LOVE CONQUERS ALL BOUNDARIES BUT TRASCENDS ALL MISFORTUNE
HOPE TRANSCENDS EACH RESTRICTION LOVE CONQUERS ALL BOUNDARIES BUT TRASCENDS EVERY MISFORTUNE
@@
HOPE TRIUMPHS AGAINST OBSTACLES LOVE CONQUERS ALL DIFFERENCES AND TRASCENDS ALL ADVERSITY
HOPE TRIUMPHS OVER CHALLENGES LOVE CONQUERS ALL DIFFERENCES AND TRASCENDS ALL MISFORTUNE
HOPE VANQUISHES ALL HINDRANCES LOVE CONQUERS ALL DIFFERENCES BUT TRASCENDS ALL MISFORTUNE
LOVE OVERCOMES EACH HINDRANCES LOVE CONQUERS ALL DIFFERENCES BUT TRASCENDS EVERY MISFORTUNE
LOVE CONQUERS ALL ADVERSITY LOVE CONQUERS ALL OBSTACLES AND TRASCENDS ALL ADVERSITY
LOVE CONQUERS ALL BOUNDARIES LOVE CONQUERS ALL OBSTACLES AND TRASCENDS ALL MISFORTUNE
LOVE CONQUERS ALL DIFFERENCES LOVE CONQUERS ALL OBSTACLES BUT TRASCENDS ALL ADVERSITY
LOVE CONQUERS ALL OBSTACLES LOVE CONQUERS ALL OBSTACLES BUT TRASCENDS EVERY ADVERSITY